From GlueXWiki
Jump to: navigation, search

May 8, 2014 FDC+CDC meeting


  1. Gas system status [1],[2] (Dave)
  2. DAQ (Beni)
  3. CDC testing (Mike)
  4. Engineering Fluorinert level (Bill)
  5. Electronics (Chris, Nick)
  6. FDC alignment FDC E-log entry 144, [3], [4], [5],[6], [7](Lubomir)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Mike, Dave, Chris, Nick, Eugene, Simon, Beni, and Lubomir (JLab).

Gas system

- The pure gas has arrived and we started purging the gas system in the gas room, including the two mixing tanks. Beni: There's a supply now to only one tank (we are still missing one MFC control panel) and the two tanks are connected at the output lines of the tanks. Beni is afraid that there will be an air pocket in the second tank and will take time to flush it. In any case we expect the control panel to arrive soon.

- The plan is to flush the gas system in the gas room over the weekend and at the beginning of the next week to include the lines in the hall bypassing the detectors. Later next week we will start flushing the detectors, as well.

- Dave has a PLC software and GUI that can be used to open/close the valves needed for the flushing procedure. Mike is working on a GUI that uses epics variables for the gas system. Dave suspects a glitch in the MFC control panel software that, however, it can be overcome for now.


- Beni: third VXS crate has been installed in the trigger rack to be used for triggering the FDC and CDC DAQ. Still needs from Sasha modules: TS, CPU and also wants to have a fast F1TDC to read the signals from the scintillators. The "horizontal" trigger gives about 80 counts per hour, but we don't know how many of these are really horizontal muons, most are expected to be coming from mini-showers above the scintillators.

- Beni's plan is to make a common FDC+CDC trigger using the "horizontal" trigger plus the "vertical" trigger from the scintillators above and below the magnet, as used by BCAL. Nick will take care of the cable connections between the scintillator electronics and the trigger rack.

- DAQ software: the readout code compiled but not tested. The sparsification code is ready and has been used in the FDC readout; before using it Beni will have to find the right timing windows.

CDC status

- Mike: 4 signal cables need to be replaced. Problems with two HV modules CAEN 1550P. Nick replaced them with spares, but one of them still shows a problem: current is high even without anything connected to that channel. No more spares, and will take few weeks for the modules to be repaired, therefore CAEN 1535P can be used. 1550P has finer current scale, 100nA/bit while 1535P has 500nA/bit.

- Curtis: with 100nA/bit scale the modules can't be used for monitoring since the dark current in the CDC is in the order of tens of nA. Same problem is with FDC and that's why we replaced the 1535P with 1550P. However, if you want to look at individual channels you can use the Bertan HV supply (with 1nA sensitivity) using a special connector that Chris made.

- Beni and Lubomir pointed out that we had problems with the FDC when the CAEN module jumped to the max hardware HV. Therefore, the HV should be disconnected from the detector for now and the hardware limits that are about 5kV by default should be lowered (using the pot on the module and telnet to read the values, the slow control system can't read them).


- There's some Fluorinert leakage, in two weeks the level on the chiller when down by about 4-5mm. Will wait for another two weeks and measure the weight of the Fluorinert needed to fill the chiller to the same line, just to have an estimation for the costs. At the same time Dave is going to borrow (or buy) a leak detector to try to find the leakage.


- Chris checked all the channels in the Beni's list with a scope using a new board (without resistors) and put the results in the xls file. Fernando will have to look at them and decide what to do next with the 3M connectors.

FDC alignment

- Lubomir made a presentation about an alignment procedure using the cosmic data from the third package full test. The only assumption is that the spacing of the wires and the rotations of the wire frames are correct. The time-to-distance function and all the corrections to the rotations and the spacing of the cathode strips are extracted from the data. All the details are given in the FDC E-log entry 144. Interestingly, one can see some gaps between the foils of the cathodes up to 200 microns, similarly to what we found from the Bill's scan of a cathode.


- No meeting next week due to the collaboration meeting.